OpenStudy (anonymous):

How do you find the slope-intercept form of 7x-3y=-12.

OpenStudy (anonymous):

Standard form: \[Ax + By = C\]Slope-Intercept form: \[y = mx + b\] Basically, all you need to do is solve for y. \[ 7x-3y=-12 \]Subtract 7x from both sides.\[-3y= -7x -12\]Divide both sides by -3. ****-7 divided by -3 would be -7/-3, but since both the numerator and denominator are negatives, we switch the fraction to a positive. Final result.\[y = 7/3 + 4\]

OpenStudy (anonymous):

@Josh45680

OpenStudy (anonymous):

How do you subtract from both sides? Why is 12 still there?

OpenStudy (anonymous):

Ok lets go through this step by step. 7x - 3y = -12 To get that into Slope-Intercept form, we need to isolate 'y'. We want to eliminate 7x from the left side of the equation. Since its a positive (7x), we subtract 7x from both sides: 7x - 7x - 3y = -12 - 7x So your left with -3y = -12 - 7x. You get it so far?

OpenStudy (anonymous):

Umm.. Sure. Thanks for your help